OpenStudy (anonymous):

4^(x + 7) = 6^(x – 1) x ≈ 0.04 x ≈ 0.05 x ≈ 19.51 x ≈ 28.35

OpenStudy (callisto):

Similar to your previous question ...again... Take log on both sides. log[4^(x + 7)] = log[6^(x – 1)] (x+7) log4 = (x-1) log6 Distribute, simplify and solve it.

OpenStudy (anonymous):

can i send u a link about this question

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what do i distribute?

OpenStudy (callisto):

I thought you called it as ''distribute''... It's actually like this: (x+7) log4 = (x-1) log6 x (log4) + 7 (log4) = x(log6) - 1(log6)